OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t^2 + 704t. After how many seconds does the projectile take to reach its maximum height?

OpenStudy (lgbasallote):

maximum height is when velocity is zero so take the derivative then equate to zero

OpenStudy (lgbasallote):

do you get that?

mathslover (mathslover):

\[\large{\frac{u^2\sin^2\theta}{2g}=H}\]

mathslover (mathslover):

Wait lemme think about the correct way about this .. I do think that equations of motion will be applied here

mathslover (mathslover):

Ok so here i go with the correct soln

mathslover (mathslover):